# LLaRA: Large Language-Recommendation Assistant

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington D.C., USAisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
  • ID-based Representation: Within the prompt, each item is represented as an ID number (Geng et al., 2022) (e.g., “14” for the movie “Titanic”) or a randomly-initialized ID token (Hua et al., 2023), as Figures 0(a) and 0(b) illustrate, respectively. Despite its simplicity, this approach leaves the textual characteristics of items (e.g., titles and descriptions) untouched, consequently underutilizing the world knowledge inherent in LLMs. Moreover, the employment of ID numbers or ID tokens might pose integration challenges with LLMs, as it does not correspond well with the natural language processing capabilities of LLMs.
@@ 2. Related Work
(a) Curriculum Prompt Tuning for Hybrid Item Representation.
(b) Architecture of SR2LLM.
@@ 4.2. Hybrid Prompt Design
Our hybrid prompt design facilitates an integration of textual metadata and ID-based item embeddings sourced from a well-trained recommender model. This design addresses the limitations of prompts that rely exclusively on either ID-based or textual data, thereby generating more accurate recommendations.
@@ 5.1. Experimental Settings
Traditional GRU4Rec 1.0000 0.3750 1.0000 0.4535 1.0000 0.2616
Caser 1.0000 0.3861 1.0000 0.4885 1.0000 0.2233
SASRec 1.0000 0.3444 1.0000 0.5282 1.0000 0.2233
LLM-based Llama2 0.4421 0.0421 0.1661 0.0361 0.3443 0.0246
GPT-4 0.9895 0.2000 0.9782 0.2039 1.0000 0.3770
MoRec 1.0000 0.2822 1.0000 0.3956 1.0000 0.1652
TALLRec 0.9263 0.3895 0.9648 0.6787 0.9836 0.4180
Ours LLaRA (GRU4Rec) 0.9684 0.4421 0.9966 0.7139 0.9836 0.4344
LLaRA (Caser) 0.9684 0.4737 0.9992 0.7072 0.9918 0.4344
LLaRA (SASRec) 0.9684 0.4421 0.9916 0.6955 1.0000 0.4508
@@ 5.2.2. Results
(a) LLaRA outperforms all baselines on all three datasets. Specifically, it achieves the highest HitRatio@1 metric of 0.4737, 0.7139 and 0.4508 on MovieLens, Steam and LastFM, respectively. This validates its effective integration of traditional sequential information with the extensive world knowledge and robust reasoning capabilities of LLMs.
@@
(d) LLaRA achieves a high validity ratio of over 95% on all datasets, illustrating the model’s instruction-following abilities when generating recommendations. It’s worth noting that all generative methods that incorporate LLMs might generate invalid answers. For instance, Llama2, which serves as the backbone LLM of LLaRA, only achieves valid ratios of 0.4421, 0.1661, and 0.3443 on the MovieLens, Steam, and LastFM datasets, respectively. Remarkably, LLaRA’s significant improvement in valid ratios can be attributed to the fact that LLaRA has been instruction-tuned on the sequential recommendation task.
@@ 5.4. Impact of Curriculum Prompt Tuning (RQ3)
Direct 0.4211 0.6921 0.4508
Two-stage 0.4316 0.6779 0.4344
LLaRA (CL) 0.4421 0.6955 0.4508

### v4

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657690isbn: 979-8-4007-0431-4/24/07ccs: Information systems Recommender systems
@@ 1. Introduction
  • ID-based Representation: Within the prompt, each item is represented as an ID number (Geng et al., 2022) (e.g., “14” for the movie “Titanic”) or a randomly-initialized ID token (Hua et al., 2023), as Figures 1(a) and 1(b) illustrate, respectively. Despite its simplicity, this approach leaves the textual characteristics of items (e.g., titles and descriptions) untouched, consequently underutilizing the world knowledge inherent in LLMs. Moreover, the employment of ID numbers or ID tokens might pose integration challenges with LLMs, as it does not correspond well with the natural language processing capabilities of LLMs.
@@ 2.2. Multi-Modal Large Language Models
(a) Curriculum Prompt Tuning for Hybrid Item Representation.
(b) Architecture of SR2LLM.
@@ 4.2. Hybrid Prompt Design
Our hybrid prompt design facilitates integration of textual metadata and ID-based item embeddings sourced from a well-trained recommender model. This design addresses the limitations of prompts that rely exclusively on either ID-based or textual data, thereby generating more accurate recommendations.
@@ 5.1. Experimental Settings
Traditional GRU4Rec 1.0000 0.3750 1.0000 0.4168 1.0000 0.2616
Caser 1.0000 0.3861 1.0000 0.4368 1.0000 0.2233
SASRec 1.0000 0.3444 1.0000 0.4010 1.0000 0.2233
LLM-based Llama2 0.4421 0.0421 0.1653 0.0135 0.3443 0.0246
GPT-4 0.9895 0.2000 0.9798 0.3626 1.0000 0.3770
MoRec 1.0000 0.2822 1.0000 0.3911 1.0000 0.1652
TALLRec 0.9263 0.3895 0.9840 0.4637 0.9836 0.4180
Ours LLaRA (GRU4Rec) 0.9684 0.4421 0.9975 0.4924 0.9836 0.4344
LLaRA (Caser) 0.9684 0.4737 0.9966 0.4874 0.9918 0.4344
LLaRA (SASRec) 0.9684 0.4421 0.9975 0.4949 1.0000 0.4508
@@ 5.2.2. Results
(a) LLaRA outperforms all baselines on all three datasets. Specifically, it achieves the highest HitRatio@1 metric of 0.4737, 0.4949 and 0.4508 on MovieLens, Steam and LastFM, respectively. This validates its effective integration of traditional sequential information with the extensive world knowledge and robust reasoning capabilities of LLMs.
@@
(d) LLaRA achieves a high validity ratio of over 95% on all datasets, illustrating the model’s instruction-following abilities when generating recommendations. It’s worth noting that all generative methods that incorporate LLMs might generate invalid answers. For instance, Llama2, which serves as the backbone LLM of LLaRA, only achieves valid ratios of 0.4421, 0.1653, and 0.3443 on the MovieLens, Steam, and LastFM datasets, respectively. Remarkably, LLaRA’s significant improvement in valid ratios can be attributed to the fact that LLaRA has been instruction-tuned on the sequential recommendation task.
@@ 5.4. Impact of Curriculum Prompt Tuning (RQ3)
Direct 0.4211 0.4899 0.4508
Two-stage 0.4316 0.4840 0.4344
LLaRA (CL) 0.4421 0.4949 0.4508